# A Modular and Multimodal Generative AI Framework for Urban Building Energy Data: Generating Synthetic Homes

### v1

@@ Abstract
Computational models have emerged as powerful tools for energy modeling research, touting scalability and quantitative results. However, these models require a plethora of data, some of which is inaccessible, expensive, or raises privacy concerns. We introduce a modular multimodal framework to produce this data from publicly accessible residential information and images using generative artificial intelligence (AI). Additionally, we provide a pipeline demonstrating this framework, and we evaluate its generative AI components. Our experiments show that our framework’s use of AI avoids common issues with generative models. Our framework produces realistic, labeled data. By reducing dependence on costly or restricted data sources, we pave a path towards more accessible and reproducible research.
@@ 1 Introduction
Consumption of electricity in the United States rose approximately from 1980 to 2009 [1] and seasonal electricity use per person has nearly doubled from 1973 to 2025 [2]. Rising energy consumption exerts increasing pressure on infrastructure capacity, operational costs, and greenhouse gas emissions, as well as costs for end-users. Accurate modeling and policy tools are therefore essential for effective planning, aiding by preventing infrastructure strain, optimizing energy costs, minimizing or mitigating climate impacts, assessing policy effectiveness, and guiding decision-making in urban centers [3, 4, 5, 6].
@@
Computational approaches to monitor energy usage and model urban centers, buildings, and other structures require large amounts of data, such as type and quality of materials, floor plans, and weather patterns to produce scalable and quantitative results [7]. Moreover, several issues arise when attempting to collect such information, including high prices, limited availability, and privacy concerns. Consequently, these issues create barriers to accessing data, hindering the development of accurate models and scaling them to larger populations or geographic areas. The inaccessibility of this data has led researchers to generate synthetic data through a variety of methods. We propose a generative Artificial Intelligence (AI) based approach to generate low-cost synthetic data that can used to develop, validate, and improve existing computational models for energy demand modeling. The proposed method enables the generation of large volumes of data for use in any geographic region where basic descriptive data on housing structures is publicly available.
@@
Current data generation efforts span a wide range of approaches, from optimizing building construction processes to the creation of digital twins. Notably, Wan et al. [8] built a generative model that improves the efficiency of energy consumption by determining the optimal layout of floor plans. Elnabawi et al. [9] enhanced their building energy models using highly detailed microclimate models that considered features such as shadow generation and seasons to generate weather data. Lee et al. [10] train a neural network that predicts building energy consumption, aiming to create a more robust model with limited real-world measurements. The network was trained with data synthesized using electricity, temperature, and humidity, among others, encoded into receiver operating characteristic (ROC) plots which visually encode the separation between classes and are commonly used to assess model discrimination [11, 12].
@@
Other researchers focus on creating and applying digital twins: virtual, dynamic, and data-driven digital replicas of real-world urban environments. Stinner et al. [13] generate digital twins of building energy systems from piping and instrumentation diagrams using convolutional networks and other algorithms. Agostinelli et al. [14] create digital twins of energy efficient buildings in residential areas from Building Information Modeling (BIM), Internet of Things (IoT), Geographical Information Systems (GIS), and AI components, with machine learning serving as the central component of the model. Francisco et al. [15] take a different approach by generating digital twins of daily, time-segmented energy benchmarks solely from smart meter data streams. Finally, Belik and Rubanenko [16] present a hybrid digital twin framework for renewable energy sources (RES) that integrates real-time data, 3D modeling, and sensitivity analysis to improve forecasting and operational efficiency, particularly in grid systems with high RES penetration.
@@
While prior work has focused on traditional simulation tools or sensor data to construct digital twins, Xu et al. [17] encourage the use of generative AI models, like GPT, because they present new opportunities in urban data generation. For instance, Dodge et al. [18] employ convolutional neural networks (CNNs) to parse floor plan images, generating structured representations using wall segmentation, object recognition, and optical character recognition. Relatedly, Zhang et al. [19] implement a pipeline that generates EnergyPlus IDF files from initial building geometry and textual building descriptions. Moreover, Xiao and Xu [20] propose a multi-agent Large Language Model (LLM) framework that processes unstructured building data and sensor logs to support automated energy optimization in smart grid contexts. Additionally, in closely related fields, generative AI has been successfully employed to perform a variety of tasks, such as predictive analysis of traffic congestion [21], real-time decision-making in urban air mobility [22], and stochastic solar irradiance forecasting for building facades [23].
@@
Liu et al. [24] present a wide-ranging review on the integration of generative AI into building energy workflows, highlighting both opportunities and challenges. These include concerns such as dataset preparation, hallucinations, and a lack of domain-specific expertise. However, the impact of such challenges varies depending on the pipeline design. Because our framework combines pretrained LLMs with simulation tools for quantitative evaluation, we avoid the need for fine-tuning or physics-aware generation, thereby reducing the impact of these concerns. We evaluate our framework’s LLM components to reduce hallucinations. Additionally, Liu emphasizes the importance of multimodal integration—where multiple data modalities (such as image, text, and tabular data) are used by a process or model. Producing high-quality structured urban data suitable for tasks such as modeling and simulation requires combining multiple complimentary processes into a single workflow [25, 26, 27]. We adopt these principles in the design of our general framework.
In the following sections, we provide an in depth view of our framework (Section 2). Then, we test our pipeline’s generative components (Section 3), ensuring the images are being processed correctly and the labels produced by the labeler are reflective of the inputs. Finally we conclude and highlight future directions from this work (Section 4).
@@ 2 Methodology
We propose a novel framework that combines multimodal real-world data with data generated using LLMs. We leverage open-access county data and generative AI to construct synthetic homes that mirror real-world counterparts and inspection notes corresponding to each. We then label this generated data using generative AI and a formula that ranks simulation results based on worst-case and best-case values. We test the Machine Learning (ML) components of our framework, extending beyond procedural evaluations. Our framework comprises of five modular components (see Figure 1), where each stage focuses on a transformation of the data, from data collection to labeling:
1.
A web scraper that collects data, images, and floor plans;
- 2.
3.
A Geographic JSON (GeoJSON) and inspection note generator using OpenAI’s GPT-4.1-mini [29];
4.
An EnergyPlus111EnergyPlus™ is a trademark of the United States Department of Energy. [30] simulation; and
5.
A labeling system that combines heuristic rules with GPT-based natural language inference.
While our pipeline is tailored to generate labeled data from images and tabular data, it is inherently modular and extensible. Any stage is easily modified or expanded for other related purposes, and different information can be generated in addition to or in place of any data produced by this pipeline. Rather than proposing a rigid set of steps, we present a flexible framework that integrates generative AI, web scraping, and building simulation, designed for rapid, low-cost data generation in building energy and other fields that traditionally suffer from data scarcity. Ultimately, the data generated by our framework can be used to prototype, design, train, and research urban energy systems.
An instance of this pipeline was implemented and executed222Our source code is available at: https://github.com/Lafayette-EshbaughSilveyra-Group/synthetic-homes. on a g3.medium flavor JetStream 2 virtual machine running Ubuntu 24.04.2 (Noble). Specifically, the system featured one NUMA node with 8 AMD EPYC-Milan processors, 30 GB of RAM, and an NVIDIA A100 SXM4 40 GB GPU (of which 10 GB was allocated). Additionally, it was fully virtualized (KVM and AMD-V). To run the pipeline, we paid $0.36 (USD) for OpenAI requests to produce data for 257 homes—about $0.0014 per home. We used the PyTorch [31] machine learning framework, and the Hugging Face Transformers library [32] to load the models we used. In the subsections that follow, we describe each pipeline component in detail.
@@ 2.1 Scraping the Data
To ensure the the synthetic data we generated reflects real-world homes, we grounded it in data collected from county datasets. Specifically, we gathered data from Northampton County, Pennsylvania, where two of the authors work and reside. To collect this data, we implemented a web scraper using Selenium WebDriver [33] and ChromeDriver [34]. For each residential property on each street in the list of streets to process, the scraper extracts key attributes, such as number of rooms, total floor area, and other metadata listed in Table 1. This data is saved to a JavaScript Object Notation (JSON) file for each home. Additionally, the scraper downloads two images: a street view photograph and a floor plan, both of which are made available on the County’s public-facing platform. An example of each is depicted in Figure 2.
@@ 2.2 Processing Images
After scraping the images and floor plans from the county portal, they are analyzed using LLaVA [28]. This step processes the images to translate them into a textual description. For example, the floor plans can be used to define the geometry of the home, and the image to determine the quality of the windows. The use of images allows our pipeline to infer important aspects of the homes that will be used to generate the GeoJSON, which is ultimately used to generate the IDF file required by EnergyPlus.
The image processing step originally was conducted using GPT. However, our experiments (see Section 3.1) displayed lack of detail in the analyses produced by that model. We pivoted to LLaVA and found that it outperformed GPT in every experiment. We hypothesize that this advantage stems from LLaVA’s architecture: it employs an early fusion design in which images are converted into visual tokens by CLIP [35] that are then prepended to the textual prompt [28]. These visual tokens are passed through the transformer layers of the LLM together with text tokens, enabling multimodal attention at every layer. It has been demonstrated that vision tokens are most critical for a model in early layers, when they fuse into text tokens [36]. A likely result of this is the reduction of the model’s sensitivity to variations in the image, which may explain its stronger performance in our experiments.
2.3 Generating GeoJSON
Once the images are processed, we utilize the descriptions given by LLaVA in a GPT prompt to generate a GeoJSON file and home inspection notes. The prompt provides the image descriptions (Section 2.2) and data that was scraped from the county portal (Section 2.1) and directs the model to take two actions:
1.
The prompt instructs the model to generate a GeoJSON file for the building, including geometry, data from the county, and five estimated performance perimeters. These estimated parameters, defined according to standard convention, include the HVAC heating and cooling coefficients of performance, r-values for the roof and wall, and the air change rate.
- 2.
2.4 Running EnergyPlus Simulations
To execute a simulation in EnergyPlus, we take the GeoJSON generated by GPT (Section 2.3) and convert it to an Input Definition File (IDF) file for use in EnergyPlus. We use a template IDF file, filled with default values that are replaced with variables from the GeoJSON data, such as the geometry, HVAC heating and cooling coefficients of performance, r-values for the wall and roof, and the air change rate. Additionally, the template has many fields with constant values, such as the simulation running period and materials, constructions, and zones used in the simulation. Once the details from the GeoJSON are copied into the template, a final IDF file is produced.
2.5 Determining Energy Efficiency of Homes
The energy efficiency of homes is measured using two components: (1) HVAC and insulation efficiency and (2) inspection notes. The first component is derived from the EnergyPlus simulation results (Section 2.4), while the second component is generated using GPT (Section 2.3). To prepare the data, we take the results from each EnergyPlus simulation and corresponding the home inspection note and label them using two scores. The first score pertains to the HVAC system, while the second score pertains to insulation. The process of computing these scores requires three steps:
1.
Rate the need of replacing or improving HVAC systems and/or insulation by using GPT to parse the home inspection notes. We refer to this score as and its value is in .
2.
Calculate a heuristic score for HVAC systems and/or insulation based on the EnergyPlus simulation results. This score is computed using Equation 1.
(1) where is the measured value of the home being scored, represents the home that is the most efficient for each retrofit category, and represents the home that is the least efficient for each retrofit category.
For example, suppose homes and consume kWh, kWh, kWh, and kWh with their HVAC systems, respectively. Then, kWh and kWh, meaning scores , scores , scores , and scores . In this scenario, and would be prioritized for HVAC retrofitting, since they offer greater potential energy savings.
This naïve formula is not the only option; it can be replaced for other formulas that score the EnergyPlus simulation results. A modified version of the logic above could incorporate additional weights which reflect size, age, or a combination of these and other variables.
3.
Create a numerical rating by normalizing the combined results from steps 1 and 2. This is calculated by taking the weighted average of and , as depicted in Equation 2.
We write the inspection notes, EnergyPlus simulation results, and labels for each example into one JSON Lines (JSONL) file.
This completes our end-to-end dataset generation process, yielding a unified dataset that combines textual and simulation-based information with consistent labeling. By design, each component of the pipeline can be adapted, replaced, or extended for use in other domains or research questions, making the framework a reusable foundation for diverse modeling and simulation tasks.
3 Experiments & Results
There is great risk that tools used within the pipeline could impact the results due to their architecture. To mitigate some of those concerns, we conducted experiments that validate the output produced by our pipeline according to the input received. Two of the main areas that showed the greatest bias in our pipeline were the image processing component from Section 2.2 and the labeler from Section 2.5. Some of these tests were executed on Lafayette College’s high-performance computing cluster, Firebird, while others were run on a local machine. In the following subsections we address each case.
3.1 Evaluating Visual Focus with Occlusion Testing
@@
To compare the focus of the GPT and LLaVA LLMs when analyzing images, we ran occlusion tests. We provided the models with twenty different images of homes: ten with roofs showing no signs of damage, and ten with roofs with significant damage. Each model was instructed to evaluate and describe the state of the roof and justify its evaluation. We first took a baseline where no part of the image was masked. Then, the image was divided into one hundred cells, a number we selected to balance effectiveness, efficiency, and cost. Each cell of the image was masked, and the model was executed on the original prompt and this new image. Textual responses were embedded into vectors using the all-MiniLM-L6-v2 SentenceTransformers model [45], a lightweight and widely used model sufficient for obtaining sentence-level numerical representations. Then, to measure the difference between each execution of the experiment, we took the cosine distance of each embedded response from the baseline. Additionally, heatmaps of these differences were assembled using matplotlib [46].
The mean difference value in roof cells and non-roof cells for each image was calculated. Figure 4 depicts the distribution of the forward occlusion results. These scatter plots are particularly telling—GPT (Figure 4(a)) evenly scatters the data points on either side of the line , indicating GPT has no preference for which area in the image is considered important. In comparison, LLaVA (Figure 4(b)) tends to produce results in the upper half of the graph. This illustrates LLaVA’s improvement over GPT plainly: more often than not, LLaVA’s response depends more on the roof than not.
@@
When we compared individual GPT and LLaVA occlusion results, we found more evidence that LLaVA performs better at focusing on key parts of the images. LLaVA is much more particular about which cells are key to construct the response; GPT produces results that look stochastic. Figure 6 displays one noteworthy example—LLaVA focuses on a more specific section of the image, primarily consisting of the home’s roof. GPT, on the other hand, behaves much more randomly. Based on these results, we selected LLaVA for use in the image processing step of our pipeline. Full experimental details and findings are available in B.
3.2 Validation of Labeling Components
As discussed in Section 2.5, the pipeline includes a labeler that quantifies the efficiency of homes. The labeler consists of two components: a labeler for the EnergyPlus simulation results and a labeler for the home inspection notes. Our goal is to achieve a labeler whose outputs are meaningful, and that balances the contributions of each component. In our study, we conducted an experiment to measure the sensitivity of each component. Given the multimodal nature of our pipeline, the first two parts of this experiment employ Ablation, with the third building on the findings of the original two.
Ablation testing is commonly used to evaluate multimodal models by focusing on one modality at a time [47]. It can reveal biases, weak model components, and hallucinations by fixing all but one of a model’s modality inputs and varying the unfixed one. Biases could impact a labeling module like ours by considering certain modalities of the data more important than others. This can have a negative effect on the overall labels assigned, as some of the data within the dataset is used less—or perhaps not used at all—to label the dataset.
We applied two ablation tests to reveal any biases between the EnergyPlus simulation results and inspection note components in our labeling model. Each test varied the efficiency of the home in one component, keeping the other component constant. Results were plotted to show the variation in the output given the efficiency of the varied component. To complement the ablation tests, we ran an additional test in which we paired every possible combination of home inspection notes and simulation results representing efficient and inefficient homes. For each prompt, we ran five trials, and the mean and standard deviation are reported. Full experimental values are provided in A. We present three versions of the labelers, iteratively mitigating bias in each.
3.2.1 Naïve Labeler
Originally, we used GPT to generate the labels, feeding it both the simulation data and the inspection notes. However, we found that this architecture was heavily skewed toward the inspection notes, such that those features disproportionally determined the label output, even when paired with simulation data (see Figure 7). This skew is especially evident when viewing the text ablation results (Figure 7(a)) and the combined variation results (Figure 8(b)). In the text ablation results, as the inspection notes move from very inefficient (HVAC 1) to very efficient (HVAC 5), there is a dramatic reduction in score. Clearly, the labeler is extremely sensitive to text. However, it is much less sensitive to simulation results. In fact, in the EnergyPlus ablation test (Figure 8(a)), we would expect mean HVAC label values to trend to decrease as the index increases, since as index increases, Coefficient of Performance (COP) also increases. Yet this trend does not occur; instead mean HVAC label values fluctuate. This bias perpetuates into the combined variation tests, which show a clear favoring of text data.
3.2.2 Heuristic Labeler
To reduce bias, we introduced the heuristic labeler. GPT processes the text and a heuristic labeler processes the simulation data, using the heuristic labeling equations described in Section 2.5. The results of this iteration—displayed in Figure 9—were an advancement from the previous. We observed the desired decreasing trend in the EnergyPlus ablation test (Figure 9(b)). At this stage, EnergyPlus results reflected the inputs; with an increase in COP, the HVAC score decreases. In the combined variation test (Figure 10(a)), we observed that the HVAC scores given were high when the inspection note described an inefficient home, and low when it described an efficient home; the heuristic labels were still lightly weighted in comparison to GPT’s inspection note labels. The current results do not produce a labeler where an input describing a home that is efficient in a given category and paired with inefficient simulation data produces a result similar to an input describing an inefficient home paired with efficient simulation data. In other words, the two middle results of any combined variation test, like the ones presented in Figure 10(a), should have a similar mean label value.
3.2.3 Weighted Heuristic Labeler
To balance the weights between text and simulation data, we employ a weighted sum. The text score is weighed at and the simulation results at . With this updated methodology, our results are more equitable (see Figure 11). The EnergyPlus ablation test (Figure 11(b)) showed a much more pronounced decreasing trend, and combined variation tests (Figures 12(a) and 12(b)) reveal that the model more equally weighs between the two modalities. For both HVAC- and insulation-related variables, in combined input variation testing, we observed that inspection notes and data describing efficient homes scored the lowest, notes and data describing inefficient homes scored the highest, and mixed-efficiency notes and data scored between the previous values. Therefore, the labeler now accurately assigns reasonable retrofits.

### v2

# Synthetic Homes: An Accessible Multimodal Pipeline for Producing Residential Building Data with Generative AI

@@ Abstract
Computational models have emerged as powerful tools for energy modeling research, touting scalability and quantitative results. However, these models require a plethora of data, some of which can be inaccessible, expensive, or can raise privacy concerns. We introduce a modular multimodal framework to produce this data from publicly accessible images and residential information using generative Artificial Intelligence (AI). Additionally, we provide a pipeline demonstrating this framework and we evaluate its generative AIcomponents. Our experiments show that our framework’s use of AIavoids common issues with generative models and produces realistic multimodal data. By reducing dependence on costly or restricted data sources, we pave a path towards more accessible research in Machine Learning (ML)and other data-driven disciplines.
@@
Novel modular framework to generate synthetic data for use in urban energy research.
Uses textual and visual analysis for generating data with LLMs for direct use in BEM.
Data generated is validated to reduce biases that result from the use of LLMs.
Includes statistical evidence that shows generated data is realistic.
Produces data while spending less than US $1 per home for GPT usage and electricity.
@@ 1 Introduction
Consumption of electricity in the United States will increase by between 2020 and 2026 [1] and seasonal electricity use per person has nearly doubled from 1973 to 2025 [2]. Rising energy consumption exerts increasing pressure on infrastructure capacity, operational costs, and greenhouse gas emissions, resulting in rising costs for governments, energy producers, and energy consumers. Accurate modeling and policy tools are therefore essential for effective planning, aiding by preventing infrastructure strain, optimizing energy costs, minimizing or mitigating climate impacts, assessing policy effectiveness, and guiding decision-making in urban centers [3, 4, 5, 6].
@@
Computational approaches to monitor energy usage and model urban centers, buildings, and other structures require large amounts of data, such as type and quality of materials, floor plans, and weather patterns to produce scalable and quantitative results [7]. However, several issues arise when attempting to collect such information, including high prices, limited availability, and privacy concerns. These issues create barriers to accessing data, hindering the development of accurate models and the scaling of models to larger populations or geographic areas. The inaccessibility of this data has led researchers to generate synthetic data through a variety of methods. We propose a generative AIbased approach to generate low-cost multimodal synthetic data consisting of simulations and home inspection notes that can be used to develop, validate, and improve existing computational models for energy demand modeling. The proposed method enables the generation of large volumes of data for use in any geographic region where basic descriptive data on housing structures is publicly available.
In the following sections, we position our work in the existing literature (Section 2) and provide an in depth view of our framework (Section 3). Then, we evaluate the realism of data generated by our pipeline (Section 4) using a dataset created to statistically model the region of our synthetic homes. Finally we conclude and highlight future directions from this work (Section 5).
2 Related Works
@@
Current data generation efforts span a wide range of approaches, from optimizing building construction processes to the creation of digital twins. For example, Wan et al. [8] built a generative model that improves the efficiency of energy consumption by determining the optimal layout of floor plans. Elnabawi et al. [9] enhanced their building energy models using highly detailed microclimate models that considered features such as shadow generation and seasons to generate weather data. Lee et al. [10] train a neural network that predicts building energy consumption, aiming to create a more robust model with limited real-world measurements. The network was trained with data synthesized using, among others, electricity, temperature, and humidity, encoded into receiver operating characteristic (ROC) plots which visually encode the separation between classes and are commonly used to assess model discrimination [11, 12].
@@
Other researchers focus on creating and applying digital twins: virtual, dynamic, and data-driven digital replicas of real-world urban environments. Stinner et al. [13] generate digital twins of building energy systems from piping and instrumentation diagrams using convolutional networks and other algorithms. Agostinelli et al. [14] create digital twins of energy efficient buildings in residential areas from Building Information Modeling (BIM), Internet of Things (IoT), Geographical Information Systems (GIS), and AIcomponents, with machine learning serving as the central module of the model. Francisco et al. [15] take a different approach by generating digital twins of daily, time-segmented energy benchmarks solely from smart meter data streams. Belik and Rubanenko [16] present a hybrid digital twin framework for renewable energy sources (RES) that integrates real-time data, 3D modeling, and sensitivity analysis to improve forecasting and operational efficiency, particularly in grid systems with high RES penetration.
@@
While prior work has focused on traditional simulation tools or sensor data to construct digital twins, Xu et al. [17] encourage the use of generative AImodels, such members of the GPTfamily, because they present new opportunities in urban data generation. For instance, Dodge et al. [18] employ convolutional neural networks (CNNs) to parse floor plan images, generating structured representations using wall segmentation, object recognition, and optical character recognition. Relatedly, Zhang et al. [19] implement a pipeline that generates EnergyPlusIDF files from initial building geometry and textual building descriptions. Moreover, Xiao and Xu [20] propose a multi-agent Large Language Model (LLM)framework that processes unstructured building data and sensor logs to support automated energy optimization in smart grid contexts. Additionally, in closely related fields, generative AIhas been successfully employed to perform a variety of tasks, such as predictive analysis of traffic congestion [21], real-time decision-making in urban air mobility [22], and stochastic solar irradiance forecasting for building facades [23].
@@
Liu et al. [24] present a wide-ranging review on the integration of generative AIinto building energy workflows, highlighting both opportunities and challenges. These include concerns such as dataset preparation, hallucinations, and a lack of domain-specific expertise. However, the impact of such challenges varies depending on the pipeline design. Because our framework combines pretrained LLMswith simulation tools for quantitative evaluation, we avoid the need for fine-tuning or physics-aware generation, thereby reducing the impact of these concerns. We evaluate our framework’s LLMcomponents to reduce hallucinations. Additionally, Liu emphasizes the importance of multimodal integration—where multiple data modalities (such as image, text, and tabular data) are used by a process or model. Producing high-quality structured urban data suitable for tasks such as modeling and simulation requires combining multiple complementary processes into a single workflow [25, 26, 27]. We adopt these principles in the design of our general framework.
@@ 3 Methodology
We propose a novel framework that combines multimodal real-world data with data generated using LLMs. We leverage open-access county data and generative AIto construct synthetic homes that mirror real-world counterparts and inspection notes corresponding to each. Home inspection notes will be used as a source to rate the homes based on their energy efficiency, alongside home images, floorplans, and simulation results. Finally, we test the LLMcomponents of our framework and the realism of the generated data. Our framework comprises four modular components, depicted in Figure 1, where each stage focuses on a transformation of the data, from data collection to simulation:
While our pipeline is tailored to generate data from images and tabular data, it is inherently modular and extensible. Any stage is easily modified or expanded for other related purposes, and different information can be generated in addition to or in place of any data produced by this pipeline. For example, a different LLMcould be employed in place of GPT, or other types of data could be considered or generated by the pipeline. Those interested in generating synthetic commercial buildings could use a modified version of this pipeline that emphasizes the characteristics and data central to their research. This opens use of our work to a larger set of individuals. Rather than proposing a rigid set of steps, we present a flexible framework that integrates generative AI, web scraping, and building simulation, designed for rapid, low-cost data generation in building energy and other fields that traditionally suffer from data scarcity. Ultimately, the data generated by our framework can be used to prototype, design, train, and research urban energy systems and models.
An instance of this pipeline was implemented and executed222Our source code is available at: https://github.com/Lafayette-EshbaughSilveyra-Group/synthetic-homes. on a g3.medium flavor JetStream 2 virtual machine running Ubuntu 24.04.2 (Noble). Specifically, the system featured one NUMA node with 8 AMD EPYC-Milan processors, 30 GB of RAM, and an NVIDIA A100 SXM4 40 GB GPU (of which 10 GB was allocated). Additionally, it was fully virtualized (KVM and AMD-V). We paid $0.36 (USD) for OpenAI requests to produce data for 258 homes—about $0.0014 per home.333Assuming a conservative upper bound of 10 CPU minutes per home at 100 watts and electricity prices at $0.15 per kilowatt-hour, local electricity costs are approximately $0.0025 per synthetic home, giving a worst case total of $0.0039 per synthetic home. To run the pipeline, we used the PyTorch [31] machine learning framework, and the Hugging Face Transformers library [32] to load the models we used. In the subsections that follow, we describe each pipeline component in detail.
@@ 3.1 Scraping the Data
To ensure the synthetic data we generated reflects real-world homes, we collected data from publicly available county datasets. For our pipeline instance, we gathered data from Northampton County, Pennsylvania, where two of the authors work and reside. To collect this data, we implemented a web scraper using Selenium WebDriver [33] and ChromeDriver [34]. For each residential property on each street in the list of streets to process, the scraper extracts key attributes, such as number of rooms, total floor area, and other metadata listed in Table 1. This data is saved to a JavaScript Object Notation (JSON)file for each home. Additionally, the scraper downloads two images: a street view photograph and a floor plan, both of which are made available on the County’s public-facing platform. An example of each is depicted in Figure 2.
@@ 3.2 Processing Images
After scraping the street view photographs and floor plans from the county portal, they are analyzed using LLaVA [28]. This step processes both files to translate them into a textual description. For example, the floor plans can be used to define the geometry of the home, and the image to determine the quality and number of windows. The use of images allows our pipeline to infer important aspects of the homes that will be used to generate the GeoJSON, which is ultimately used to generate the Input Definition File (IDF)file required by EnergyPlus.
3.2.1 Evaluating Visual Focus with Occlusion Testing
@@
During the development of our pipeline, we identified two candidate LLMsto employ for image processing. To determine the best model for this task, we compared the focus of the GPTand LLaVALLMswhen analyzing images using occlusion tests. We provided the models with twenty different images of homes: ten with roofs showing no signs of damage, and ten with roofs with significant damage. Each model was instructed to evaluate and describe the state of the roof and justify its evaluation. We first took a baseline where no part of the image was masked. Then, the image was divided into one hundred cells, a number we selected to balance effectiveness, efficiency, and cost. Each cell of the image was masked, and the model was executed on the original prompt and this new image. Textual responses were embedded into vectors using the all-MiniLM-L6-v2 SentenceTransformers model [37], a lightweight and widely used model sufficient for obtaining sentence-level numerical representations. Then, to measure the semantic difference between each execution of the experiment, we took the cosine distance of each embedded response from the baseline. Additionally, heatmaps of these differences were assembled using matplotlib [38].
The mean difference value in roof cells and non-roof cells for each image was calculated. Figure 3 depicts the distribution of the forward occlusion results. These scatter plots are particularly telling—GPT(Figure 3(a)) evenly scatters the data points on either side of the line , indicating it has no preference for which area in the image is considered important. In comparison, LLaVA(Figure 3(b)) tends to produce results in the upper half of the graph. This illustrates LLaVA’s improvement over GPTplainly: more often than not, LLaVA’s response depends more on the roof than other parts of the image.
@@
When we compared individual GPT and LLaVAocclusion results, we found more evidence that LLaVAperforms better at focusing on key parts of the images. LLaVAis much more particular about which cells are key to construct the response; GPT produces results that look stochastic. Figure 5 displays one noteworthy example—LLaVAfocuses on a more specific section of the image, primarily consisting of the home’s roof. GPT, on the other hand, behaves much more randomly. Each value in the cell is the semantic difference, and the background tint of the cell represents the scaled semantic difference, where the brightest background represents the largest difference and the darkest represents the smallest. In Figure 5(a), the semantic difference is generally larger on the roof than on other parts of the image. A small semantic difference is a signal that a given portion of the image is less important to the model in order to formulate a response. In Figure 5(b), the majority of cells are moderately important to the model in formulating a response—even cells containing walls or grass. Comparing the two, LLaVAdemonstrates better focus on the more important parts of the image. Full occlusion experimental details, hypotheses, and findings are available in A.
3.3 Generating GeoJSON
Once the images are processed, we utilize the descriptions given by LLaVAin a GPTprompt (see Appendix B) to generate a GeoJSONfile and home inspection notes. The prompt provides the image descriptions (Section 3.2) and data that was scraped from the county portal (Section 3.1) and directs the model to take two actions:
1.
The prompt instructs GPTto generate a GeoJSONfile for the building, including geometry, data from the county, and five estimated performance parameters. These estimated parameters, defined according to standard convention, include the HVAC heating and cooling coefficients of performance, r-values for the roof and wall, and the air change rate.
2.
The prompt directs GPTto write a short inspection note, focused on energy-related observations such as insulation, HVAC type/age, visible windows, and any inferred upgrades. This inspection note is placed in the GeoJSONfile.
3.4 Running EnergyPlus Simulations
To execute a simulation in EnergyPlus, we take the GeoJSONgenerated by GPT(Section 3.3) and convert it to an IDFfile for use in EnergyPlus. We use a template IDFfile, filled with default values that are replaced with variables from the GeoJSONdata, such as the geometry, HVAC heating and cooling coefficients of performance, r-values for the wall and roof, and the air change rate. Additionally, the template has many fields with constant values, such as the simulation running period and materials, constructions, and zones used in the simulation. Once the details from the GeoJSONare copied into the template, a final IDFfile is produced.
We write the inspection notes and EnergyPlus simulation results for each example into one JSON Lines (JSONL)file. This completes our end-to-end dataset generation process, yielding a unified dataset that combines textual and simulation-based information.
4 Validating Realism of Pipeline Generated Data
Following prior work on synthetic data evaluation [42, 43], we adopt the definition of realism as the overlap of synthetic values with the – range of a trusted reference on the home characteristics being generated. To measure the realism of our generated data, we compare it with ResStock [44, 45], a dataset that contains realistic samples of homes in different regions of the United States as IDFfiles. The analysis consists of performing statistical analyses comparing the distribution of values in our synthetic homes to those in ResStock. We only use ResStock data from climate zone 4A, the region in which the synthetic homes are collected in our example, and we only compare the four pre-selected variables—Wall R-Value, Roof R-Value, Cooling COP, Heating COP—that are not static among all synthetic homes. The number of synthetic homes is while the number of climate zone 4A ResStock samples is .
Table 2 reports the realism for the four variables across all synthetic homes by displaying the percentage of values that overlap with the – range of ResStock. We found overlap greater than for all variables and greater than for three variables. This overlap is a strong indicator of realism for our generated data. Additionally, the table includes the mean and median of the four variables for all synthetic homes and the interquartile range (IQR) of the four variables for all ResStock homes. These statistics are included for informational purposes and are not considered when determining realism.
| Variable | Overlap (%) | ResStock IQR (Q1–Q3) | Synthetic Mean | Synthetic Median |
| Wall R-Value | 93.41 | 0.704 (1.94–2.64) | 2.79 | 2.29 |
| Roof R-Value | 100.0 | 3.35 (3.35–6.69) | 5.04 | 5.28 |
| Cooling COP | 66.28 | 0.70 (3.1–3.8) | 2.99 | 3.00 |
| Heating COP | 100.0 | 0.20 (0.8–1.0) | 0.84 | 0.85 |
The results above and the tests of the generative components of our framework demonstrate a level of confidence above previous work such as Zhang et al. [19] and Xiao and Xu [20]. These works and others either accept the output of the LLMsor use procedural evaluation techniques.
Occlusion offers stronger empirical support that the image processing component of our pipeline—which is LLM-based—performs as expected. We additionally demonstrated statistically that properties of our generated homes are realistically distributed according to ResStock. Our results directly address and mitigate concerns about hallucination and unreliability raised in prior work [24]. It is fair to be skeptical of LLMsand other MLmodels—but with our results, we clearly demonstrate our pipeline is not as susceptible to some of the issues Liu et al. [24] raise.